# MuMustango: Toward Controllable Text-to-Music Generation

### v2

@@ Abstract
[…]
@@ 3.2 MuNet
$A_{\tau}=\text{MHA}_{\theta_{\tau}}(Q=z_{n},K/V=\text{FLAN-T5}(\tau))$
@@
$\epsilon_{\theta}^{(n)}(z_{n},\mathcal{C}):=\text{UNet}_{\theta}(A_{c})$ (2)
@@
MuNet consists of a UNet Ronneberger et al. (2015)—consisting of multiple downsampling, middle, and upsampling blocks—and multiple conditioning cross-attention blocks. We use two encoders, $\boldsymbol{Enc^{b}}$ and $\boldsymbol{Enc^{c}}$, to encode the beat and chord features which leverage both the state-of-the-art Fundamental Music Embedding (FME) as well as an onset-and-beat-based positional encoding Guo et al. (2023) which we name Music Positional Encoding (MPE). These ensure the musical features are properly captured and several fundamental music properties (e.g., intervals between pitches are translational invariant) are preserved.
@@ Appendix C Performance of the Predictors
Prompt: “This folk song features a female voice singing the main melody. This is accompanied by a tabla playing the percussion. A guitar strums chords. For most parts of the song, only one chord is played. At the last bar, a different chord is played. This song has minimal instruments. This song has a story-telling mood. This song can be played in a village scene in an Indian movie. The chord sequence is Bbm, Ab. The beat is 3. The tempo of this song is Allegro. The key of this song is Bb minor. Without control sentences in italics (TestA): chords predicted: ["G", "C", "G", "C", "G", "C"], chords predicted time: [0.46, 1.21, 3.25, 5.48, 7.24, 8.92]. chords extracted from audio: ["G6", "C", "G", "C", "G", "Cmaj7"], chords time extracted from audio: [0.46, 1.58, 3.07, 5.94, 7.62, 9.66] With control sentences in italics (TestB): chords predicted: ["Bbm", "Ab"], chords predicted time: [0.46, 7.24], chords extracted from audio: ["F#maj7", "Ab"], chords time extracted from audio: [0.46, 7.43].
Prompt: “A female singer sings this bluesy melody. The song is medium tempo with minimal guitar accompaniment and no other instrumentation. The song’s medium tempo is very emotional and passionate. The song is a modern pop hit but with poor audio quality. The key of this song is G minor. The time signature is 3/4. This song goes at 168.0 beats per minute. The chord progression in this song is Am7, G7, Cm, G, A7. Without control sentences in italics (TestA): chords predicted: ["C#m7", "C#m7", "C#m7", "C#m7", "C#m7"], chords predicted time: [0.46, 3.25, 6.32, 8.17, 9.29], chords extracted from audio: ["F#", "C#m", "F#m", "C#m7"], chords time extracted from audio: [0.46, 1.21, 4.55, 5.39] With control sentences in italics (TestB): chords predicted: ["Am7", "G7", "Cm", "G", "A7"], chords predicted time: [0.46, 1.67, 3.53, 5.48, 8.92], chords extracted from audio: ["Am", "G", "C", "Gmaj7", "A6", "Gmaj7"], chords time extracted from audio: [0.46, 1.67, 3.72, 5.94, 8.73, 9.85]
@@ Appendix H ChatGPT Rephrasing
"I have a song for which the caption is the following:" *caption* "I have made some changes to the audio file which are optionally described towards the end of the caption. Can you rephrase the caption more naturally in a single paragraph using all the musical terms provided above? You should generate only the caption and nothing else. Do not use the word modification in your generation. The length of the new caption should be no more than eight sentences."
Examples of rephrasing: Original This folk song features a male voice singing the main melody in an emotional mood. This is accompanied by an accordion playing fills in the background. A violin plays a droning melody. There is no percussion in this song. This song can be played at a Central Asian classical concert. The chord sequence is B, Dmaj7, B, Gbm. This song goes at 95.0 beats per minute. The beat counts to 2. This song is in the key of B minor.
Rephrased A male voice sings the main melody with emotional intensity in this Central Asian folk song. The accordion fills the background while a violin provides a droning melody. There is no percussion in this piece, which can be played at a classical concert in Central Asia. With a chord sequence of B, Dmaj7, B, Gbm, the song has a beat count of 2 at a tempo of 95.0 beats per minute. The song is in the key of B minor.
Original This folk song features a male voice singing the main melody in an emotional mood. This is accompanied by an accordion playing fills in the background. A violin plays a droning melody. There is no percussion in this song. This song can be played at a Central Asian classical concert. The key is C minor. The chord progression in this song is C, D#maj7, C, Gm.
Rephrased This emotional folk song, perfect for a Central Asian classical concert, showcases a male voice singing the main melody accompanied by a droning violin and accordion fills in the background. With no percussion present, the key of C minor sets the tone, and the chord progression follows suit with C, D#maj7, C, Gm.
@@ Appendix I Control-Sentence Templates to Enhance the Prompts
This piece is an instrumental reggae song that is very chill and slow. There is no singer. It is relaxing to hear the groove with the bass guitar. The song includes reggae electric guitar, horn, and percussion like bongos. The keyboard provides lush chords. The time signature is 4/4. The chord progression is G, F, C.
2
@@
This instrumental blues song goes very slow at a bpm of 50. You can hear the bass, harmonica and guitar grooving. The harmonica plays a solo over the harmonious guitar and bass.
3
@@
This classical piece is a waltz played by a string quartet. It includes two violins, a viola, and a cello, the beat counts to 3. It sounds elegant, and has a strong first beat. It has a natural and danceable rhythm. The mood is romantic. The chord progression is Em, Am, D, G.
4
@@
African drums are playing a complex rhythm while a male vocalist chants a ritual. The atmosphere is mesmerizing. The complex drumming pattern is a mesmerizing blend of syncopation, polyrhythms, and intricate patterns. It takes place somewhere in the wilderness, or in an indigenous village.
5
@@
This rock piece with guitars and drums is loud but fades out later on and becomes softer. It sounds powerful yet melancholic. It is instrumental only. A bass guitar provides a steady beat, enhancing the groove and energy of the song.
6
@@
A single bass instrument is playing a running baseline. It has a jazzy feeling to it and sounds mellow. This could be played in a jazz club. The tempo is 120 bpm.
7
@@
This is a hip hop song. It has two rappers taking turns, one female and one male. An electronic synth melody sample in the background keeps on looping. We can hear electronic beats and sometimes record-scratching sound effects.
8
@@
A smooth jazz song with saxophone, drums and guitar with a chord progression of Dm7, G7, Cmaj7. The song is relaxed and slow. There are no vocals, it is instrumental only. The saxophone produces a velvety tone that delivers an emotive melody.
9
@@
A piano plays a soothing popular instrumental song that could serve as background music in a restaurant. There is only piano playing, no other instruments. There is a piano melody with background piano chords of Am, Fmaj7, Cmaj7, and G. The tempo is unhurried. The melody is gentle and soothing, evoking a sense of nostalgia and comfort.
10
@@
Indian folk music with a sitar and female vocals. It evokes a sense of zen and elevation. A sitar player begins with a gentle and melodic introduction, plucking the strings with precision and emotion. There are rhythmic beats of traditional hand percussion instruments, such as the tabla. It could be played at a cultural festival to showcase Indian culture.
11
@@
This is a melodic and energetic rock ballad with a male vocalist. It has a country vibe and is of alternative or popfolk genre. The electric and acoustic guitars and the bass create the background, while the drums give a regular beat. The singer’s voice is complemented by a piano.
12
@@
This is a slow classical piece with violins and pianos. It has a film score feel and is instrumental only. The orchestration is soft, with strings and flutes.
13
@@
This fast and energetic rock song is performed by a male singer. The genre is alternative or punk rock. The background is formed by a guitar, an electric guitar, bass, and drums. There is also a synthesizer.
14
@@
This is a slow and ambient instrumental piece with a soundscape that feels like space. The atmosphere is meditative and relaxing but with a certain darkness to it. The genre is electronic soundtrack, and the music is completely instrumental with a synthesizer, bass, and drums forming the background. This song goes at 167.0 beats per minute.
15
@@
This is an instrumental piece with Indian and classical elements. The sitar, violin, and flute play prominent roles in creating a meditative and relaxing mood. The percussion and guitar provide a background rhythm to this world and jazz fusion.
@@
An instrumental blues melody played by a lead guitar and a strumming acoustic guitar. The acoustic guitarist’s strumming keeps the rhythm steady. The chord sequence is G7, F7, C7, G7. This song goes at 100 beats per minute.
2
@@
An instrumental blues melody played by a lead guitar and a strumming acoustic guitar. The acoustic guitarist’s strumming keeps the rhythm steady. The chord sequence is Dm, Am, Em. This song goes at 60 beats per minute.
3
@@
A piano plays a popular melody over the chords of Am, Fmaj7, Cmaj7, G. There is only piano playing, no other instruments or voice. The tempo is Adagio.
4
@@
A piano plays a popular melody over the chords of Gm, Bb, Eb. There is only piano playing, no other instruments or voice. The tempo is Vivace.
5
@@
This is an intense and loud punk song with guitars and drums. It is instrumental only. It is very energetic and powerful. The thunderous beats of the drummer provide a pounding rhythm. A guitar solo melody emerges from the chaotic background of the chords. The chord progression is A, D, E. The tempo of the song is 160 bpm.
6
@@
This is an intense and loud punk song with guitars and drums. It is instrumental only. It is very energetic and powerful. The thunderous beats of the drummer provide a pounding rhythm. A guitar solo melody emerges from the chaotic background of the chords.The chord progression is C, B, A, G. The tempo of the song is 100 bpm.
7
@@
A slow paced jazz song played by a saxophone, piano, guitar and drums follows a chord progression of Em7b5, A7, Dm7. The pianist produces delicate harmonies and subtle embellishments. The drummer provides a brushed rhythm. The guitar strums softly, while the saxophone plays a solo over the chords. This song goes at 80 beats per minute.
8
@@
A slow paced jazz song played by a saxophone, piano, guitar and drums follows a chord progression of B7, G7, E7, C7. The drummer provides a brushed rhythm. The guitar strums softly, while the saxophone plays a solo over the chords. This song goes at 115 beats per minute.
9
@@
This is a techno piece with drums and beats and a leading melody. A synth plays chords. The music kicks off with a powerful and relentless drumbeat. Over the pounding beats, a leading melody emerges. It has strong danceability and can be played in a club. The tempo is 120 bpm. The chords played by the synth are Am, Cm, Dm, Gm.
10
@@
This is a techno piece with drums and beats and a leading melody. A synth plays chords. The music kicks off with a powerful and relentless drumbeat. Over the pounding beats, a leading melody emerges. It has strong danceability and can be played in a club. The tempo is 160 bpm. The chords played by the synth are C, F, G.
11
@@
A horn and a bass guitar groove to a reggae tune. The combination of the horn section’s catchy melodies and the buoyant bassline creates an irresistible groove. The bassline is bouncy and lively. The song is played at the pace of Adagio. An electric keyboard plays the chords Am, Dm, G, C.
12
@@
A horn and a bass guitar groove to a reggae tune. The combination of the horn section’s catchy melodies and the buoyant bassline creates an irresistible groove. The bassline is bouncy and lively. The song is played at the pace of Moderato. An electric keyboard plays the chords E, B, A.
13
@@
This is a metal song with a guitar, drums and bass guitar. The bassist, wielding a solid-bodied bass guitar, adds depth and power to the sonic landscape. The drummer commands a massive drum kit. With a relentless force, they pound out thunderous rhythms, driving the music forward. As the song begins, the guitar roars to life, delivering a series of distorted chords. It follows the chords of Em, C, G, D. The tempo is 120 bpm.
14
@@
This is a metal song with a guitar, drums and bass guitar. The bassist, wielding a solid-bodied bass guitar, adds depth and power to the sonic landscape. The drummer commands a massive drum kit. With a relentless force, they pound out thunderous rhythms, driving the music forward. As the song begins, the guitar roars to life, delivering a series of distorted chords. It follows the chords of A, F#m, D, E. The tempo is 170 bpm.
15
@@
A man sings a captivating folk song while strumming chords on an acoustic guitar. This fits a campfire evening happening. The chord progression is G, C, D, G. The tempo is 100 beats per minute.
16
@@
A man sings a captivating folk song while strumming chords on an acoustic guitar. This fits a campfire evening happening. The chord progression is Am, Em, Dm, Am. The tempo is 70 beats per minute.
17
@@
This is a classical music piece played by a string trio. The instruments involved are violin, viola, and cello. The violin plays the lead melody. The cello’s soulful and melodic contributions add depth and gravitas to the performance. The time signature is ¾. The tempo of this song is Presto. The chord sequence is E, C#m, A, B.
18
@@
This is a classical music piece played by a string trio. The instruments involved are violin, viola, and cello. The violin plays the lead melody. The cello’s soulful and melodic contributions add depth and gravitas to the performance. The time signature is 4/4. The tempo of this song is Andante. The chord sequence is Am, Dm, E7, Am.
19
@@
This is a pop song with a female singer singing the leading melody and synthesizers looping samples as background. These loops provide the song’s electronic foundation, creating a rich and layered sonic landscape. The charismatic female singer has a dynamic and emotive voice. The tempo is Moderato. The chord sequence is C, G, Am, F.
20
@@
This is a pop song with a female singer singing the leading melody and synthesizers looping samples as background. These loops provide the song’s electronic foundation, creating a rich and layered sonic landscape. The charismatic female singer has a dynamic and emotive voice. The tempo is Presto. The key is A minor and the chord sequences are Am, Dm, E.

### v3

@@ Abstract
Jan Melechovsky1thanks:    Co-first authors. Both authors contributed equally. , Zixun Guo211footnotemark: 1 , Deepanway Ghosal1, Navonil Majumder1, Dorien Herremans1thanks:    Both authors contributed equally and led this project., Soujanya Poria122footnotemark: 2 1 Singapore University of Technology and Design, Singapore 2 Queen Mary University of London, UK
@@ 3.2 MuNet
$U^{(1)}=z_{n}$
$A_{\tau}=\text{MHA}_{\theta_{\tau}}(Q=U^{(l)},K/V=\text{FLAN-T5}(\tau))$
@@
$U^{(l+1)}=\text{UNet}^{(l)}_{\theta}(A_{c})$
$\epsilon_{\theta}^{(n)}(z_{n},\mathcal{C}):=U^{(L+1)}$ (2)
@@
MuNet consists of a UNet Ronneberger et al. (2015)—consisting of in total $L$ downsampling, middle, and upsampling blocks—and multiple conditioning cross-attention blocks. We use two encoders, $\boldsymbol{Enc^{b}}$ and $\boldsymbol{Enc^{c}}$, to encode the beat and chord features which leverage both the state-of-the-art Fundamental Music Embedding (FME) as well as an onset-and-beat-based positional encoding Guo et al. (2023) which we name Music Positional Encoding (MPE). These ensure the musical features are properly captured and several fundamental music properties (e.g., intervals between pitches are translational invariant) are preserved.
@@ Appendix C Performance of the Predictors
{mdframed}
[backgroundcolor=blue!10] Prompt: “This folk song features a female voice singing the main melody. This is accompanied by a tabla playing the percussion. A guitar strums chords. For most parts of the song, only one chord is played. At the last bar, a different chord is played. This song has minimal instruments. This song has a story-telling mood. This song can be played in a village scene in an Indian movie. The chord sequence is Bbm, Ab. The beat is 3. The tempo of this song is Allegro. The key of this song is Bb minor.
Without control sentences in italics (TestA): chords predicted: ["G", "C", "G", "C", "G", "C"], chords predicted time: [0.46, 1.21, 3.25, 5.48, 7.24, 8.92]. chords extracted from audio: ["G6", "C", "G", "C", "G", "Cmaj7"], chords time extracted from audio: [0.46, 1.58, 3.07, 5.94, 7.62, 9.66]
With control sentences in italics (TestB): chords predicted: ["Bbm", "Ab"], chords predicted time: [0.46, 7.24], chords extracted from audio: ["F#maj7", "Ab"], chords time extracted from audio: [0.46, 7.43].
{mdframed}
[backgroundcolor=blue!10] Prompt: “A female singer sings this bluesy melody. The song is medium tempo with minimal guitar accompaniment and no other instrumentation. The song’s medium tempo is very emotional and passionate. The song is a modern pop hit but with poor audio quality. The key of this song is G minor. The time signature is 3/4. This song goes at 168.0 beats per minute. The chord progression in this song is Am7, G7, Cm, G, A7.
Without control sentences in italics (TestA): chords predicted: ["C#m7", "C#m7", "C#m7", "C#m7", "C#m7"], chords predicted time: [0.46, 3.25, 6.32, 8.17, 9.29], chords extracted from audio: ["F#", "C#m", "F#m", "C#m7"], chords time extracted from audio: [0.46, 1.21, 4.55, 5.39]
With control sentences in italics (TestB): chords predicted: ["Am7", "G7", "Cm", "G", "A7"], chords predicted time: [0.46, 1.67, 3.53, 5.48, 8.92], chords extracted from audio: ["Am", "G", "C", "Gmaj7", "A6", "Gmaj7"], chords time extracted from audio: [0.46, 1.67, 3.72, 5.94, 8.73, 9.85]
@@ Appendix H ChatGPT Rephrasing
{mdframed}
[backgroundcolor=blue!10] "I have a song for which the caption is the following:" *caption* "I have made some changes to the audio file which are optionally described towards the end of the caption. Can you rephrase the caption more naturally in a single paragraph using all the musical terms provided above? You should generate only the caption and nothing else. Do not use the word modification in your generation. The length of the new caption should be no more than eight sentences."
Examples of rephrasing: {mdframed}[backgroundcolor=yellow!10] Original
This folk song features a male voice singing the main melody in an emotional mood. This is accompanied by an accordion playing fills in the background. A violin plays a droning melody. There is no percussion in this song. This song can be played at a Central Asian classical concert. The chord sequence is B, Dmaj7, B, Gbm. This song goes at 95.0 beats per minute. The beat counts to 2. This song is in the key of B minor.
{mdframed}
[backgroundcolor=green!10] Rephrased
A male voice sings the main melody with emotional intensity in this Central Asian folk song. The accordion fills the background while a violin provides a droning melody. There is no percussion in this piece, which can be played at a classical concert in Central Asia. With a chord sequence of B, Dmaj7, B, Gbm, the song has a beat count of 2 at a tempo of 95.0 beats per minute. The song is in the key of B minor.
{mdframed}
[backgroundcolor=yellow!10] Original
This folk song features a male voice singing the main melody in an emotional mood. This is accompanied by an accordion playing fills in the background. A violin plays a droning melody. There is no percussion in this song. This song can be played at a Central Asian classical concert. The key is C minor. The chord progression in this song is C, D#maj7, C, Gm.
{mdframed}
[backgroundcolor=green!10] Rephrased
This emotional folk song, perfect for a Central Asian classical concert, showcases a male voice singing the main melody accompanied by a droning violin and accordion fills in the background. With no percussion present, the key of C minor sets the tone, and the chord progression follows suit with C, D#maj7, C, Gm.
@@ Appendix I Control-Sentence Templates to Enhance the Prompts
1 This piece is an instrumental reggae song that is very chill and slow. There is no singer. It is relaxing to hear the groove with the bass guitar. The song includes reggae electric guitar, horn, and percussion like bongos. The keyboard provides lush chords. The time signature is 4/4. The chord progression is G, F, C.
@@
2 This instrumental blues song goes very slow at a bpm of 50. You can hear the bass, harmonica and guitar grooving. The harmonica plays a solo over the harmonious guitar and bass.
@@
3 This classical piece is a waltz played by a string quartet. It includes two violins, a viola, and a cello, the beat counts to 3. It sounds elegant, and has a strong first beat. It has a natural and danceable rhythm. The mood is romantic. The chord progression is Em, Am, D, G.
@@
4 African drums are playing a complex rhythm while a male vocalist chants a ritual. The atmosphere is mesmerizing. The complex drumming pattern is a mesmerizing blend of syncopation, polyrhythms, and intricate patterns. It takes place somewhere in the wilderness, or in an indigenous village.
@@
5 This rock piece with guitars and drums is loud but fades out later on and becomes softer. It sounds powerful yet melancholic. It is instrumental only. A bass guitar provides a steady beat, enhancing the groove and energy of the song.
@@
6 A single bass instrument is playing a running baseline. It has a jazzy feeling to it and sounds mellow. This could be played in a jazz club. The tempo is 120 bpm.
@@
7 This is a hip hop song. It has two rappers taking turns, one female and one male. An electronic synth melody sample in the background keeps on looping. We can hear electronic beats and sometimes record-scratching sound effects.
@@
8 A smooth jazz song with saxophone, drums and guitar with a chord progression of Dm7, G7, Cmaj7. The song is relaxed and slow. There are no vocals, it is instrumental only. The saxophone produces a velvety tone that delivers an emotive melody.
@@
9 A piano plays a soothing popular instrumental song that could serve as background music in a restaurant. There is only piano playing, no other instruments. There is a piano melody with background piano chords of Am, Fmaj7, Cmaj7, and G. The tempo is unhurried. The melody is gentle and soothing, evoking a sense of nostalgia and comfort.
@@
10 Indian folk music with a sitar and female vocals. It evokes a sense of zen and elevation. A sitar player begins with a gentle and melodic introduction, plucking the strings with precision and emotion. There are rhythmic beats of traditional hand percussion instruments, such as the tabla. It could be played at a cultural festival to showcase Indian culture.
@@
11 This is a melodic and energetic rock ballad with a male vocalist. It has a country vibe and is of alternative or popfolk genre. The electric and acoustic guitars and the bass create the background, while the drums give a regular beat. The singer’s voice is complemented by a piano.
@@
12 This is a slow classical piece with violins and pianos. It has a film score feel and is instrumental only. The orchestration is soft, with strings and flutes.
@@
13 This fast and energetic rock song is performed by a male singer. The genre is alternative or punk rock. The background is formed by a guitar, an electric guitar, bass, and drums. There is also a synthesizer.
@@
14 This is a slow and ambient instrumental piece with a soundscape that feels like space. The atmosphere is meditative and relaxing but with a certain darkness to it. The genre is electronic soundtrack, and the music is completely instrumental with a synthesizer, bass, and drums forming the background. This song goes at 167.0 beats per minute.
@@
15 This is an instrumental piece with Indian and classical elements. The sitar, violin, and flute play prominent roles in creating a meditative and relaxing mood. The percussion and guitar provide a background rhythm to this world and jazz fusion.
@@
1 An instrumental blues melody played by a lead guitar and a strumming acoustic guitar. The acoustic guitarist’s strumming keeps the rhythm steady. The chord sequence is G7, F7, C7, G7. This song goes at 100 beats per minute.
@@
2 An instrumental blues melody played by a lead guitar and a strumming acoustic guitar. The acoustic guitarist’s strumming keeps the rhythm steady. The chord sequence is Dm, Am, Em. This song goes at 60 beats per minute.
@@
3 A piano plays a popular melody over the chords of Am, Fmaj7, Cmaj7, G. There is only piano playing, no other instruments or voice. The tempo is Adagio.
@@
4 A piano plays a popular melody over the chords of Gm, Bb, Eb. There is only piano playing, no other instruments or voice. The tempo is Vivace.
@@
5 This is an intense and loud punk song with guitars and drums. It is instrumental only. It is very energetic and powerful. The thunderous beats of the drummer provide a pounding rhythm. A guitar solo melody emerges from the chaotic background of the chords. The chord progression is A, D, E. The tempo of the song is 160 bpm.
@@
6 This is an intense and loud punk song with guitars and drums. It is instrumental only. It is very energetic and powerful. The thunderous beats of the drummer provide a pounding rhythm. A guitar solo melody emerges from the chaotic background of the chords.The chord progression is C, B, A, G. The tempo of the song is 100 bpm.
@@
7 A slow paced jazz song played by a saxophone, piano, guitar and drums follows a chord progression of Em7b5, A7, Dm7. The pianist produces delicate harmonies and subtle embellishments. The drummer provides a brushed rhythm. The guitar strums softly, while the saxophone plays a solo over the chords. This song goes at 80 beats per minute.
@@
8 A slow paced jazz song played by a saxophone, piano, guitar and drums follows a chord progression of B7, G7, E7, C7. The drummer provides a brushed rhythm. The guitar strums softly, while the saxophone plays a solo over the chords. This song goes at 115 beats per minute.
@@
9 This is a techno piece with drums and beats and a leading melody. A synth plays chords. The music kicks off with a powerful and relentless drumbeat. Over the pounding beats, a leading melody emerges. It has strong danceability and can be played in a club. The tempo is 120 bpm. The chords played by the synth are Am, Cm, Dm, Gm.
@@
10 This is a techno piece with drums and beats and a leading melody. A synth plays chords. The music kicks off with a powerful and relentless drumbeat. Over the pounding beats, a leading melody emerges. It has strong danceability and can be played in a club. The tempo is 160 bpm. The chords played by the synth are C, F, G.
@@
11 A horn and a bass guitar groove to a reggae tune. The combination of the horn section’s catchy melodies and the buoyant bassline creates an irresistible groove. The bassline is bouncy and lively. The song is played at the pace of Adagio. An electric keyboard plays the chords Am, Dm, G, C.
@@
12 A horn and a bass guitar groove to a reggae tune. The combination of the horn section’s catchy melodies and the buoyant bassline creates an irresistible groove. The bassline is bouncy and lively. The song is played at the pace of Moderato. An electric keyboard plays the chords E, B, A.
@@
13 This is a metal song with a guitar, drums and bass guitar. The bassist, wielding a solid-bodied bass guitar, adds depth and power to the sonic landscape. The drummer commands a massive drum kit. With a relentless force, they pound out thunderous rhythms, driving the music forward. As the song begins, the guitar roars to life, delivering a series of distorted chords. It follows the chords of Em, C, G, D. The tempo is 120 bpm.
@@
14 This is a metal song with a guitar, drums and bass guitar. The bassist, wielding a solid-bodied bass guitar, adds depth and power to the sonic landscape. The drummer commands a massive drum kit. With a relentless force, they pound out thunderous rhythms, driving the music forward. As the song begins, the guitar roars to life, delivering a series of distorted chords. It follows the chords of A, F#m, D, E. The tempo is 170 bpm.
@@
15 A man sings a captivating folk song while strumming chords on an acoustic guitar. This fits a campfire evening happening. The chord progression is G, C, D, G. The tempo is 100 beats per minute.
@@
16 A man sings a captivating folk song while strumming chords on an acoustic guitar. This fits a campfire evening happening. The chord progression is Am, Em, Dm, Am. The tempo is 70 beats per minute.
@@
17 This is a classical music piece played by a string trio. The instruments involved are violin, viola, and cello. The violin plays the lead melody. The cello’s soulful and melodic contributions add depth and gravitas to the performance. The time signature is ¾. The tempo of this song is Presto. The chord sequence is E, C#m, A, B.
@@
18 This is a classical music piece played by a string trio. The instruments involved are violin, viola, and cello. The violin plays the lead melody. The cello’s soulful and melodic contributions add depth and gravitas to the performance. The time signature is 4/4. The tempo of this song is Andante. The chord sequence is Am, Dm, E7, Am.
@@
19 This is a pop song with a female singer singing the leading melody and synthesizers looping samples as background. These loops provide the song’s electronic foundation, creating a rich and layered sonic landscape. The charismatic female singer has a dynamic and emotive voice. The tempo is Moderato. The chord sequence is C, G, Am, F.
@@
20 This is a pop song with a female singer singing the leading melody and synthesizers looping samples as background. These loops provide the song’s electronic foundation, creating a rich and layered sonic landscape. The charismatic female singer has a dynamic and emotive voice. The tempo is Presto. The key is A minor and the chord sequences are Am, Dm, E.